Aaron saw it was 86° Fahrenheit,F, in Australia in January. Aaron used the formula C=59(F−32)

C=59(F−32)to find the temperature in degrees Celsius,C. What temperature, in degrees Celsius, did Aaron find?

Use the on-screen keyboard to type your answer in the box below.


1
Expert's answer
2021-01-19T03:50:23-0500

C=5/9(F−32)

C= 5/9(86-32)

C = 5/9(54) = 30°C.